# Discovering universal temperature regulation dynamics in animals

### v2

@@ 6.4 Model selection
To perform model selection on Arctic ground squirrel body temperature data, we used the data found in 9903_M_Toolik_temp.csv in the supplement of [9]. We opted to use a segment of the Arctic ground squirrel body temperature data set that was six spikes long. Specifically, we used the data from index 10811 to index 13747 as training data and from index 13748 to index 16573 as validation data. We applied min-max scaling to the training and validation data. We note that using different data sets may produce different model selection results (see Discussion for further information on modeling choices and limitations).
@@ 6.6 Parameter estimation on selected models
7 Supplement
7.1 Figures
Figure 6: Left: The eigenvalue spectra of the Hessian of the cost function for Model 6 at the best fit exhibits sloppiness (see [52]). Right: Eigenvector contributions from each model parameter (see [52]). As expected, $\theta_{0}$ and $y(0)$ define the stiffest directions.
@@ 7.3 Nondimensionalization and assumptions on model parameters
We make the assumption $\gamma\approx\xi$. During parameter estimation, we fix $\theta_{4}$ to make the models identifiable ($\theta_{4}=9.99995$, the value of Model 8AB that came out of model selection with $\theta_{0}=0$). After reoptimization, the numerical values of the nondimensionalized parameters are $\gamma=0.9930155091$, $\xi=0.9776338555$, $\iota=0.00002991498959$, and $\kappa=0.2621331463$. From this, we can see that $\gamma\approx\xi$. Under this assumption, we can factor the system, which yields the model we analyze:
@@ 7.4 Modified model functional forms and parameter values
The Arctic ground squirrel model
In Figure 5D, we simulate Eqs. (10)-(11) using the parameter values $\gamma=1.1,$ $\kappa=0.1,$ $\iota=1$ x 10-4, $\hat{x}_{0}=1$ x 10-4, and $\hat{y}_{0}=1$ x 10-4.
@@ 7.4.3 The black bear model
Acknowledgments

### v4

@@ 4.4 Model selection
To perform model selection on Arctic ground squirrel body temperature data, we used the data found in 9903_M_Toolik_temp.csv in the supplement of [9]. We opted to use a segment of the Arctic ground squirrel body temperature data set that was six spikes long. Specifically, we used the data from index 10811 to index 13747 as training data and from index 13748 to index 16573 as validation data (see Figures 2A and 7). Including more data was infeasible as doing so would increase the run time of the algorithm, which in some cases took up to 48 hours on an HPC cluster. We applied min-max scaling to the training and validation data. We note that using different data sets may produce different model selection results (see Discussion for further information on modeling choices and limitations).
@@ 4.6 Parameter estimation on selected models
4.7 Code
Code used for this analysis can be found here https://github.com/cefitzg/heterothermy.git.
5 Appendix
5.1 Local stability analysis
The Arctic ground squirrel thermophysiology model we discovered (Eqs. 3-4) can also describe seasonal transitions between summer homeothermy and the hibernation thermoregulatory regime. In the summer, the Arctic ground squirrel is homeothermic, with minor fluctuations induced by activity and sleep around a physiological fixed point near 37 C. During hibernation, the organism cannot remain solely at 37 C or the low state (near or below 0 $\circ$ C), but must oscillate between the two states to survive. Therefore, any physiological fixed points corresponding to high-temperature or low-temperature states must be unstable (or non-attractive). Mathematically, we expect the stability of the model’s high-temperature fixed point to transition from stable to unstable in the fall and back to stable in the spring. The model’s low-temperature fixed point should always be unstable, as staying at a low temperature without oscillations at any time of year would be fatal for the organism. To confirm this, we perform a stability analysis of our model.
$\{\text{dimensionless body temperature}\}\penalty 10000\ \penalty 10000\ \dot{\hat{x}}$ $=\hat{x}^{2}-\hat{x}^{3}-\gamma\hat{x}\hat{y}+\xi\hat{x}^{2}\hat{y}+\iota,\hat{x}(0)=\hat{x}_{0},$ (8)
$\{\text{dimensionless unknown physiological state}\}\penalty 10000\ \penalty 10000\ \dot{\hat{y}}$ $=\kappa\left(\hat{x}^{2}-\hat{y}^{2}\right),\hat{y}(0)=\hat{y}_{0}.$ (9)
$\{\text{dimensionless body temperature}\}\penalty 10000\ \penalty 10000\ \dot{\hat{x}}$ $=\hat{x}\left(\hat{x}-\gamma\hat{y}\right)\left(1-\hat{x}\right)+\iota,\hat{x}(0)=\hat{x}_{0},$ (10)
$\{\text{dimensionless unknown physiological state}\}\penalty 10000\ \penalty 10000\ \dot{\hat{y}}$ $=\kappa\left(\hat{x}^{2}-\hat{y}^{2}\right),\hat{y}(0)=\hat{y}_{0}.$ (11)
The model parameters of Eqs. (10)-(11) has interpretable physiological meaning. $\gamma$ describes the seasonal coupling between body temperature and the unknown physiological state. The parameters $\iota$ and $\kappa$ influence the period of oscillations during hibernation. Eqs. (10)-(11) has three fixed points: $(\hat{x},\hat{y})=(0,0)$, (1,1)(1,-1), and $(1,1)$ when $\iota=0.$ [^6] It is reasonable to take $\iota=0$ to examine fixed points sufficiently far away from the origin, as the higher order terms will locally dominate $\iota$, as $\iota<<1$, but this is not reasonable near the origin where $\iota$ dominates. So under the $\iota=0$ approximation, we cannot comment on the stability of the fixed point at the origin, as our assumption does not hold. Additionally, note that the fixed point $(\hat{x},\hat{y})=(1,-1)$ is unstable for all $\gamma$ and $\kappa$. A linear stability analysis of Eqs. (10)-(11) when $\iota=0$ shows that $(\hat{x},\hat{y})=(1,1)$, which corresponds to the high-temperature state, is stable for $0<\gamma<1$ and is an unstable saddle for $\gamma>1$. For $\gamma=1$, linear stability analysis is indeterminate. Biologically, the thermoregulatory transitions between homeothermy and the hibernation thermoregulatory regime can be understood as seasonal variation in $\gamma$, perhaps driven by the circannual clock, which corresponds mathematically to a seasonal change of stability of the fixed point $(\hat{x},\hat{y})=(1,1)$, the high-temperature fixed point. While the Arctic ground squirrel is homeothermic during the summer ($\gamma<1$), the thermophysiological system stays near $(\hat{x},\hat{y})=(1,1)$ with minor fluctuations. During hibernation ($\gamma>1$), $(\hat{x},\hat{y})=(1,1)$ is unstable, and the system snaps to the limit cycle, which describes the the hibernation thermoregulatory regime regime. Thus, the stability of the point near (1,1) is controlled by $\gamma$ in a seasonal manner as expected. This seasonal transition and change of stability are shown in Figure 4 for the same model with a different choice of nondimensionalization.
$1<\gamma<2(1+\kappa)$ (12)
5.2 Supporting Figures
Figure 7: Zoom in of data in Figure 2. Data: Body temperature recorded from a free-ranging Arctic ground squirrel [9]. Training data is shown in blue. Validation data is shown in orange.
Figure 8: Left: The eigenvalue spectra of the Hessian of the cost function for Model 6 at the best fit exhibit sloppiness (see [57]). Right: Eigenvector contributions from each model parameter (see [57]). As expected, $\theta_{0}$ and $y(0)$ define the stiffest directions.
@@ 5.4 Nondimensionalization and assumptions on model parameters
We make the assumption $\gamma\approx\xi$. During parameter estimation, we fix $\theta_{4}$ to make the models identifiable ($\theta_{4}=9.99995$, the value of Model 8AB that came out of model selection with $\theta_{0}=0$). After reoptimization, the numerical values of the nondimensionalized parameters are $\gamma=0.9930155091$, $\xi=0.9776338555$, $\iota=0.00002991498959$, and $\kappa=0.2621331463$. From this, we can see that $\gamma\approx\xi$. Under this assumption, we can factor the system, which yields the model we analyze in Section 5.1.
@@ 5.5 Modified model functional forms and parameter values
5.5.1 The Arctic ground squirrel model
In Figures 4 B and C and 5 D and E, we simulate Eqs. (5)-(6) using the parameter values $\omega=0.99,$ $\rho=0.1,$ $\iota=1$ x 10-4, $\tilde{x}_{0}=0.99$, and $\tilde{y}_{0}=0.99$. We take
$\nu(t)=\begin{cases}0.2(\text{tanh}(0.005(\tau-3,500)))+0.79&\text{if }0\leq\tau\leq 12,000,\\ 0.2(1-\text{tanh}(0.005(\tau-20,500)))+0.79&\text{if }12,000<\tau\leq 24,000.\end{cases}$
@@ 6.2 Funding
7 Author contributions statement
C.E.F. and N.M.M. designed research; C.E.F. and A.J.M. performed research; C.E.F and A.J.M. analyzed data; C.E.F. wrote the paper; A.J.M. and N.M.M. edited the paper.